I have problem in this question So plz solve this asap !!!!!!!


If p and q are the roots of the equation x2 - bx + c = 0, then what is the equation if the roots are (pq + p + q) and (pq - p - q)?


 

Grade:

1 Answers

vijay narang
15 years ago

I solve this question here ...this question is simple !!!

explanation:

In the given quadratic equation x2 - bx + c = 0,
The sum of the roots p + q = b --- (1)
And the product of the roots pq = c --- (2)

We have to formulate a quadratic equation whose roots are (pq + p + q) and (pq - p - q).

The sum of the two roots = pq + p +q + pq - p - q = 2pq
But from eqn (2), we know that pq = c
Therefore, the sum of the roots = 2c

The product of the roots = (pq + p + q)(pq - p - q)= (pq)2 - (p+q)2
From equation (1) and (2), we know that pq = c and p + q = b
Therefore, the product of the roots = c2 - b2

We know the sum of the roots and the product of the roots.
Therefore, the quadratic equation is x2 - (sum of the roots)x + product of the roots = 0

=> x2 - 2cx + c2 - b2 = 0